Tater and Pepper Corp. reported free cash flows for 2018 of $49.1 million and investment in operating capital of $32.1 million. Tater and Pepper incurred $14.6 million in depreciation expense and paid $30.9 million in taxes on EBIT in 2018. Calculate Tater and Pepper’s 2018 EBIT.
Answer #1
FCF = EBIT+Depreciation-Taxes-Net capital expenditure-Changes in NWC
EBIT = FCF-(Depreciation-Taxes-Net capital expenditure-Changes in NWC) =
= 49.10-(14.6-30.9-32.1) = $      97.50